# Galaxy Zoo: Cosmic Dawn – morphological classifications for over 41 000 galaxies in the Euclid Deep Field North from the Hawaii Two-0 Cosmic Dawn survey

### v1

@@ 2.3 Cutout creation and upload
Figure 3: Example cutout images shown to volunteers on Zooniverse as part of GZCD. Clockwise from top left, the subjects have angular effective radii of 1.106, 1.815, 2.483 and 1.441 arcseconds, apparent HSC-$i$ band magnitudes of 19.870, 19.354, 17.356 and 19.150, and photometric redshifts of 0.768, 0.406, 0.301 and 0.398.
As the H20 survey observations were ongoing at the time of this Galaxy Zoo iteration, the subject upload process consisted of two main subsets: one batch of subjects at project launch, and a second batch later once more observations had been taken and pre-processed by the H20 pipeline. The first subset contained 16 671 PNG subjects from the first two square degrees observed by the H20 survey, after limiting the sample according to the subject selection criteria above: these were to be fully classified by at least 40 volunteers without the aid of machine learning. The second subset consisted of 30 676 JPEG subjects from the next four square degrees of the survey: these were to be classified by fewer volunteers with the aid of Galaxy Zoo’s deep learning model, Zoobot, in an active learning cycle - see Section 3.4. From initial Zoobot tests, it was decided that the subjects should instead be in JPEG format for faster Zoobot runtimes due to reduced file sizes.
@@ 2.4 Metadata
Each subject image uploaded to Zooniverse was accompanied by metadata, which can prove useful to researchers and to dedicated volunteers for whom this data increases their engagement with the subjects. Initially, this primarily consisted of a numerical identifier and name unique to that subject, alongside model parameters from the H20 pipeline for the central object in that image: the angular effective radius (or major axis measured at detection for point-like or failed sources), the apparent HSC-$g$, -$r$, and -$i$ magnitudes, and the median photometric redshift.
@@ 3.1 Decision tree changes
Firstly, the deep HSC imaging increased the confusion noise of subjects, which in rare cases caused the H20 pipeline to treat multiple nearby galaxies as a single source. In these cases, incorrect angular scaling was thus applied to the cutout images, appearing zoomed-out with central galaxies far too small to be classifiable. In order for these subjects to be used for improving the H20 team’s reduction pipeline, the ‘Star or Artifact’ response in the Galaxy Zoo decision tree was replaced with ‘Star, Artifact, or Bad Zoom’ (also referred to as ‘Problem’ elsewhere in this work), and selecting this response now presented volunteers with another question rather than moving on to the next subject, asking volunteers which of the three it was: ‘Star’, ‘Non-star Artifact’, or ‘Bad Image Zoom’.
@@ 3.2 Tagging extremely red objects
The Zooniverse provides a built-in method for volunteers to assign keywords (‘tags’) to subjects they find particularly interesting or possess features not mentioned in the decision tree. After classifying a subject, volunteers can choose to move to the project’s Talk boards where they can leave comments about the subject, including any keywords they deem appropriate - there is no set list of keywords, so as to allow for anything unusual to be labelled. For example, subjects in Galaxy Zoo might be tagged as ‘#disturbed’, ‘#lsb’, ‘#star-forming’, and so on, now also with an emphasis on tagging gravitational lenses with ‘#lens’ following Section 3.1.
@@
As such, volunteers were tasked with tagging these rare objects with either ‘#bright-red’ or ‘#dim-red’, through instructions provided in the Galaxy Zoo tutorial, help guide, and Talk Board. As with the decision tree, volunteers were asked to focus only on the central object in order to avoid duplication of classifications (as they may have appeared as central galaxies in other subjects) or the tagging of red stars (which had likely already been filtered out of being central objects by the H20 pipeline). However, to account for any potential errors with the H20 pipeline, volunteers were allowed to tag subjects in this way if they contained EROs that were overlapping or merging with the central galaxy, or acting as a lensing arc. Additionally, to ensure completeness but avoid confusion, it was recommended that any EROs notably away from the centre could still be tagged using ‘#bright-red-companion’ or ‘#dim-red-companion’.
@@ 3.4 Zoobot active learning
For this, Zoobot used an EfficientNetB0 architecture (Tan & Le, 2019) finetuned from the ‘GZ Evo’ data set (training on previous Galaxy Zoo iterations: GZ2, Hubble, CANDELS, DECaLS, DESI, Rings; Walmsley et al., 2024). The model outputs predictions for the fractions of volunteers who would select (‘vote’ for) each response to each decision tree question across decision trees from all Galaxy Zoo iterations, except for the ‘rare features’ question due to it uniquely allowing multiple responses to be selected. Specifically, the model predicts Dirichlet distributions for the possible fractions, from which we calculate the most likely (mean) fraction as well as the uncertainty (upper and lower 90 per cent confidence intervals) on the possible fractions.
With the start of classifications for the second subset of 30 676 subjects, Zoobot was deployed alongside volunteers in an active learning cycle. First, subjects were passed through Zoobot, and a retirement threshold was set such that a subject would be retired (i.e. prevented from receiving any further volunteer classifications) if Zoobot predicted with 90 per cent confidence that fewer than 20 per cent of volunteers (vote fraction < 0.2) would select the ‘Features or Disk’ response to the first question, regardless of predicted responses for the other decision tree questions. The remaining subjects (i.e. galaxies likely to have features) would then be shown to volunteers for classification. Each week, all volunteer classifications collected to date – including the new classifications from the previous week – were used to retrain Zoobot, which was then applied to the subjects once again, and so on. Zoobot is applied to all unretired subjects, and so can ‘change its mind’ (i.e. move galaxies between the active featured set and the hidden non-featured set) as performance improves.
@@
Initial testing suggested this active learning would retire $\sim$77 per cent of galaxies early and speed up classification by a factor of $\sim$5. Later evaluation showed that Zoobot allowed the majority of subjects to be confidently retired with less than 40 volunteer classifications: the model did indeed retire 77.8 per cent of galaxies early, and 528 of these subjects did not receive any volunteer classifications as Zoobot was confident enough in its initial classifications to immediately retire them before they could be seen by volunteers. Additionally, the majority of subjects retired early were smooth galaxies, as expected: of the subjects that received less than five volunteer votes before retirement, 62 per cent were voted as ‘Smooth’ by the majority of volunteers (95 per cent by Zoobot for the 528 Zoobot-only subjects), compared to only 3 per cent (0 per cent) and 12 per cent (2 per cent) for ‘Features or Disk’ and ‘Star, Artifact, or Bad Zoom’, respectively. Through confident classifications from the model leading to early subject retirement, the average rate of classifications increased by a factor $\sim$2.4 for subjects in the second subset, with the 30676 subjects retired over 83 days compared to the 16671 subjects of the first subset that were retired over 109 days. It should be noted that this is a lower estimate for the speed-up that Zoobot active learning can provide, given the higher numbers of volunteers taking part during the first subject subset. To account for this, the speed-up can alternatively be estimated to be a factor of $\sim$2.9 given that subjects in the second subset only required a mean average of 14 volunteer classifications before retirement instead of 40. Additionally, the 77.8 per cent subjects retired early by Zoobot were skewed towards a mean average of only 6 volunteer classifications (median of 4), for which the speed-up was therefore a factor of $\sim$7.
@@ 3.5.1 Volunteer weightings and redshift debiasing
In some previous iterations of Galaxy Zoo, post-processing may be applied to the ‘raw’ vote fractions in order to improve results. One method is through weighting each volunteer’s responses according to how well the volunteer generally agrees with the consensus. These weighted vote fractions are calculated by iteratively re-weighting each user based on their general agreement with other users, although the effect on performance is typically minor (e.g. Willett et al., 2017). Additionally, given the limited resolution of the images shown to volunteers, vote fractions inevitably reflect the appearance of the galaxy, not its true underlying morphology, such that higher redshift subjects are more likely to be classified as ‘Smooth’ rather than ‘Features or Disk’. Hence in previous works, attempts have been made to ‘de-bias’ the vote fractions, estimating how subjects would have been classified at lower redshift and hence their true morphology rather than that presented to volunteers (Hart et al., 2016). A more comprehensive explanation and reasoning for Galaxy Zoo weightings and debiasing efforts can be found in Section 4.3 of Walmsley et al. (2022a), which recommends only using debiased classifications to consider populations of galaxies rather than individuals. In this work, no volunteer weighting or debiasing was applied to the results, for the reasons we will expand below.
@@
where $\overline{k}$ is the mean average consistency of that volunteers’ votes compared to those of other volunteers. However, for this work it was found that only 0.8 per cent of logged in volunteers would have a weight less than 1.0 following $\overline{k}$ values below 0.6. Previous works (Simmons et al., 2017; Willett et al., 2017; Walmsley et al., 2022a) have highlighted how the main change volunteer weighting brings is to eliminate the small percentage of volunteers who each have hundreds of classifications containing a high proportion of responses labelling subjects as ‘Star or Artifact’. This is thought to be due to previous Galaxy Zoo iterations having such a response to the first question lead directly to the next subject, allowing these volunteers to skip through many subjects quickly by performing mostly inaccurate classifications, whether that be to view more subjects or to artificially increase their classification count. However, this Cosmic Dawn iteration expands the decision tree so that the updated ‘Star, Artifact or Bad Zoom’ response now leads to further questions, a by-product of which was found to be the elimination of this behaviour: no cases of this were found in the aggregated classifications. Based on this, alongside the minor number of volunteers that would be affected by weighting, it was deemed unnecessary for the volunteers’ vote fractions to be weighted.
For debiasing vote fractions, it should be noted that the current debiasing method assumes galaxy morphological properties do not evolve significantly for those of similar intrinsic brightness and physical size. While some previous iterations of Galaxy Zoo have involved sets of subjects sharing similar, and often low, redshifts, for which this assumption approximately holds (Walmsley et al., 2022a), the Cosmic Dawn iteration’s subject set covers a much wider redshift range (primarily $0<z<1.0$, as well as many subjects at higher redshift). Additionally, photometric redshift estimation from the H20 pipeline failed for approximately one fifth of the sources in the subject set. For these reasons, no redshift-based debiasing was applied to the results, with the final vote fractions remaining ‘raw’, i.e. entirely based on the visual morphology present in the images shown to volunteers. As such, classifications of objects at higher redshifts will be more adversely affected due to lower resolution, with them being more likely to be classified as non-featured (for more information on the effect of this biasing, see Walmsley et al., 2023b).
@@ 3.5.2 Galaxy Zoo Mobile gravitational lenses
Throughout the project, in addition to the ‘rare features’ question, volunteers were asked to tag any gravitational lenses they found on the Talk boards: given the rarity of such objects, we wished to explore all possible methods for compiling as complete a sample as possible, especially given that the ‘rare-features’ question is shown last to volunteers. As the project neared completion, another method was explored: in addition to the desktop version of Zooniverse, there is also a Zooniverse app for mobile devices, intended for simple and often binary ‘Yes/No’ workflows that are compatible with the mobile app’s ‘swipe left/right’ functionality. Hence, as a follow-up to the project, a workflow for finding gravitational lenses through the mobile app was launched alongside other workflows as part of a reboot of Galaxy Zoo Mobile. It consisted of 22 587 subjects, which were simply the set of galaxies that had not been fully classified in the desktop application, and removing those predicted by Zoobot to be stars, artifacts or images with ‘Bad Image Zoom’ with a 90 per cent lower bound greater than 0.5. The workflow asked volunteers if each subject contained a gravitational lens or not, and used a retirement threshold of ten volunteer classifications given the far simpler, binary ‘Yes/No’ workflow. To improve speed and volunteer experience, an additional retirement rule was introduced such that a subject would be retired if it had six or more votes and every one of those votes was for the ‘no’ response.
@@ 4.1.1 Volunteer & classification number counts
GZCD ran for over six months from 21 October 2022 to 03 May 2023 and consisted of 47 347 subjects. 3 990 911 classifications were made by tens of thousands of volunteers: 74 per cent by the 10 274 logged-in volunteers along with 26 per cent from an unknown number of volunteers who made classifications without logging in. The weekly number of total classifications made by volunteers are provided in Figure 5. The peak at the end of November corresponds to an update put out on the Galaxy Zoo Talk boards, in particular highlighting the release of the H20 Image Viewer (see Section 2.4) and, correspondingly, the inclusion of image coordinates in the Galaxy Zoo subject metadata to view the subjects. Unfortunately, the press releases (see Section 3.3) at the end of 2022 did little to increase volunteer engagement, perhaps due to a limited ability to reach would-be volunteers compared to, say, the post on the Galaxy Zoo Blog at project launch. The rise in classifications towards the end of the project was likely due to volunteers becoming more engaged as the ‘percentage complete’ bar nears completion in addition to seeing more interesting subjects as a result of Zoobot active learning.
@@
The first subject set of 16 671 subjects was fully classified by at least 40 volunteers per subject. As discussed in Section 3.4, with the introduction of Zoobot active learning to the second subject set, 528 of the subjects were immediately and confidently retired by Zoobot without any volunteer classifications; the remaining 30 148 subjects each received at least one volunteer classification, of which 19 686 (65 per cent) had fewer than ten volunteer responses before being retired by Zoobot. Figure 6 shows the number of votes that each subject received from volunteers. For subjects with less than 40 votes, there is a skew towards low counts, due to Zoobot retiring many subjects quickly and more over time as it trains through active learning. The threshold for subject retirement is made obvious by the peak at 40 votes, which, in addition to subjects in the first subject set each receiving at least 40 votes, is amplified by 22.6 per cent of subjects in the second subject set also receiving at least 40 votes: for these cases, if Zoobot could not confidently classify a subject after being seen by a small number of volunteers, it was unlikely to ever retire it early, leading to those subjects hitting (or exceeding) the retirement threshold. There are many subjects that received slightly more than 40 votes due to subjects with just under 40 votes receiving multiple volunteers almost simultaneously, which cause them to surpass 40 votes before the automated retirement can be enacted.
Figure 6 also shows another peak in the distribution at higher classification counts, with 189 subjects (0.4 per cent of data set) receiving 66-109 votes each, with a median of 87 votes. Of these subjects, 100 have randomly distributed H20 IDs: when a subject set is completed (with all subjects having reached one or more retirement criteria), Zooniverse continues to show a sample of 100 subjects, which are still classifiable but marked by a ‘Finished!’ tag, which aids in maintaining volunteer engagement before new subjects become available. The other 89 subjects correspond to the final subject IDs in the first subject set, which all received a significant fraction of their votes ($\sim$48-93 per cent) towards the end of that set’s classification period. Why these subjects received so many classifications is uncertain, but likely due to an unknown bug with the Zooniverse platform during the course of the project runtime: the subjects do not correlate to any particular object and are not duplicate uploads, nor are they erroneously duplicated entries in the classification results, and the vast majority of the votes come from logged-in volunteers so cannot be due to bot accounts repeatedly classifying the same subjects.

### v2

@@ 2.3 Cutout creation and upload
Figure 3: Example cutout images shown to volunteers on Zooniverse as part of GZCD. Clockwise from top left, the galaxies have angular effective radii of 1.106, 1.815, 2.483 and 1.441 arcseconds, apparent HSC-$i$ band magnitudes of 19.870, 19.354, 17.356 and 19.150, and photometric redshifts of 0.768, 0.406, 0.301 and 0.398.
As the H20 survey observations were ongoing at the time of this Galaxy Zoo iteration, the subject upload process consisted of two main subsets: one batch at project launch, and a second batch later once more observations had been taken and pre-processed by the H20 pipeline. The first subset contained 16 671 PNG subjects from the first two square degrees observed by the H20 survey, after limiting the sample according to the selection criteria above: these were to be fully classified by at least 40 volunteers without the aid of machine learning. The second subset consisted of 30 676 JPEG subjects from the next four square degrees of the survey: these were to be classified by fewer volunteers with the aid of Galaxy Zoo’s deep learning model, Zoobot, in an active learning cycle - see Section 3.4. From initial Zoobot tests, it was decided that the subjects should instead be in JPEG format for faster Zoobot runtimes due to reduced file sizes.
@@ 2.4 Metadata
Each subject image uploaded to Zooniverse was accompanied by metadata, which can prove useful to researchers and to dedicated volunteers for whom this data increases their engagement with the project. Initially, this primarily consisted of a numerical identifier and name unique to that subject, alongside model parameters from the H20 pipeline for the central object in that image: the angular effective radius (or major axis measured at detection for point-like or failed sources), the apparent HSC-$g$, -$r$, and -$i$ magnitudes, and the median photometric redshift.
@@ 3.1 Decision tree changes
Firstly, the deep HSC imaging increased the confusion noise of the images, which in rare cases caused the H20 pipeline to treat multiple nearby galaxies as a single source. In these cases, incorrect angular scaling was thus applied to the cutout images, appearing zoomed-out with central galaxies far too small to be classifiable. In order for these subjects to be used for improving the H20 team’s reduction pipeline, the ‘Star or Artifact’ response in the Galaxy Zoo decision tree was replaced with ‘Star, Artifact, or Bad Zoom’ (also referred to as ‘Problem’ elsewhere in this work), and selecting this response now presented volunteers with another question rather than moving on to the next subject, asking volunteers which of the three it was: ‘Star’, ‘Non-star Artifact’, or ‘Bad Image Zoom’.
@@ 3.2 Tagging extremely red objects
The Zooniverse provides a built-in method for volunteers to assign keywords (‘tags’) to subjects they find particularly interesting or possess features not mentioned in the decision tree. After classifying a subject, volunteers can choose to move to the project’s Talk boards where they can leave comments about it, including any keywords they deem appropriate - there is no set list of keywords, so as to allow for anything unusual to be labelled. For example, subjects in Galaxy Zoo might be tagged as ‘#disturbed’, ‘#lsb’, ‘#star-forming’, and so on, now also with an emphasis on tagging gravitational lenses with ‘#lens’ following Section 3.1.
@@
As such, volunteers were tasked with tagging these rare objects with either ‘#bright-red’ or ‘#dim-red’, through instructions provided in the Galaxy Zoo tutorial, help guide, and Talk Board. As with the decision tree, volunteers were asked to focus only on the central object in order to avoid duplication of classifications (as they may have appeared as central galaxies in other cutouts) or the tagging of red stars (which had likely already been filtered out of being central objects by the H20 pipeline). However, to account for any potential errors with the H20 pipeline, volunteers were allowed to tag subjects in this way if they contained EROs that were overlapping or merging with the central galaxy, or acting as a lensing arc. Additionally, to ensure completeness but avoid confusion, it was recommended that any EROs notably away from the centre could still be tagged using ‘#bright-red-companion’ or ‘#dim-red-companion’.
@@ 3.4 Zoobot active learning
For this, Zoobot used an EfficientNetB0 architecture (Tan & Le, 2019) finetuned from the ‘GZ Evo’ data set (Walmsley et al., 2024), which consisted of subjects from the following previous Galaxy Zoo iterations: Galaxy Zoo 2 (GZ2), GZ Hubble, GZ CANDELS, GZ DECaLS, GZ DESI and GZ Rings, where CANDELS, DECaLS and DESI correspond to the Cosmic Assembly Near-infrared Deep Extragalactic Legacy Survey, Dark Energy Camera Legacy Survey and the Dark Energy Spectroscopic Instrument, respectively. The model outputs predictions for the fractions of volunteers who would select (‘vote’ for) each response to each decision tree question across decision trees from all Galaxy Zoo iterations, except for the ‘rare features’ question due to it uniquely allowing multiple responses to be selected. Specifically, the model predicts Dirichlet distributions for the possible fractions, from which we calculate the most likely (mean) fraction as well as the uncertainty (upper and lower 90 per cent confidence intervals) on the possible fractions.
With the start of classifications for the second subset of 30 676 subjects, Zoobot was deployed alongside volunteers in an active learning cycle. First, the images were passed through Zoobot, and a retirement threshold was set such that a subject would be retired (i.e. prevented from receiving any further volunteer classifications) if Zoobot predicted with 90 per cent confidence that fewer than 20 per cent of volunteers (vote fraction < 0.2) would select the ‘Features or Disk’ response to the first question, regardless of predicted responses for the other decision tree questions. The remaining subjects (i.e. galaxies likely to have features) would then be shown to volunteers for classification. Each week, all volunteer classifications collected to date – including the new classifications from the previous week – were used to retrain Zoobot, which was then applied to the images once again, and so on. Zoobot is applied to all unretired subjects, and so can ‘change its mind’ (i.e. move galaxies between the active featured set and the hidden non-featured set) as performance improves.
@@
Initial testing suggested this active learning would retire $\sim$77 per cent of galaxies early and speed up classification by a factor of $\sim$5. Later evaluation showed that Zoobot allowed the majority of subjects to be confidently retired with less than 40 volunteer classifications: the model did indeed retire 77.8 per cent of galaxies early, and 528 of these subjects did not receive any volunteer classifications as Zoobot was confident enough in its initial classifications to immediately retire them before they could be seen by volunteers. Additionally, the majority of subjects retired early were smooth galaxies, as expected: of the subjects that received less than five volunteer votes before retirement, 62 per cent were voted as ‘Smooth’ by the majority of volunteers (95 per cent by Zoobot for the 528 Zoobot-only subjects), compared to only 3 per cent (0 per cent) and 12 per cent (2 per cent) for ‘Features or Disk’ and ‘Problem’, respectively. Through confident classifications from the model leading to early subject retirement, the average rate of classifications increased by a factor $\sim$2.4 for the second subset, with the 30676 subjects retired over 83 days compared to the 16671 subjects of the first subset that were retired over 109 days. It should be noted that this is a lower estimate for the speed-up that Zoobot active learning can provide, given the higher numbers of volunteers taking part during the first subset. To account for this, the speed-up can alternatively be estimated to be a factor of $\sim$2.9 given that subjects in the second subset only required a mean average of 14 volunteer classifications before retirement instead of 40. Additionally, the 77.8 per cent retired early by Zoobot were skewed towards a mean average of only 6 volunteer classifications (median of 4), for which the speed-up was therefore a factor of $\sim$7.
@@ 3.5.1 Volunteer weightings and redshift debiasing
In some previous iterations of Galaxy Zoo, post-processing may be applied to the ‘raw’ vote fractions in order to improve results. One method is through weighting each volunteer’s responses according to how well the volunteer generally agrees with the consensus. These weighted vote fractions are calculated by iteratively re-weighting each user based on their general agreement with other users, although the effect on performance is typically minor (e.g. Willett et al., 2017). Additionally, given the limited resolution of the images shown to volunteers, vote fractions inevitably reflect the appearance of the galaxy, not its true underlying morphology, such that higher redshift objects are more likely to be classified as ‘Smooth’ rather than ‘Features or Disk’. Hence in previous works, attempts have been made to ‘de-bias’ the vote fractions, estimating how objects would have been classified at lower redshift and hence their true morphology rather than that presented to volunteers (Hart et al., 2016). A more comprehensive explanation and reasoning for Galaxy Zoo weightings and debiasing efforts can be found in Section 4.3 of Walmsley et al. (2022a), which recommends only using debiased classifications to consider populations of galaxies rather than individuals. In this work, no volunteer weighting or debiasing was applied to the results, for the reasons we will expand below.
@@
where $\overline{k}$ is the mean average consistency of that volunteers’ votes compared to those of other volunteers. However, for this work it was found that only 0.8 per cent of logged in volunteers would have a weight less than 1.0 following $\overline{k}$ values below 0.6. Previous works (Simmons et al., 2017; Willett et al., 2017; Walmsley et al., 2022a) have highlighted how the main change volunteer weighting brings is to eliminate the small percentage of volunteers who each have hundreds of classifications containing a high proportion of responses labelling objects as ‘Star or Artifact’. This is thought to be due to previous Galaxy Zoo iterations having such a response to the first question lead directly to the next subject, allowing these volunteers to skip through many subjects quickly by performing mostly inaccurate classifications, whether that be to view more subjects or to artificially increase their classification count. However, this Cosmic Dawn iteration expands the decision tree so that the updated ‘Star, Artifact or Bad Zoom’ response now leads to further questions, a by-product of which was found to be the elimination of this behaviour: no cases of this were found in the aggregated classifications. Based on this, alongside the minor number of volunteers that would be affected by weighting, it was deemed unnecessary for the volunteers’ vote fractions to be weighted.
For debiasing vote fractions, it should be noted that the current debiasing method assumes galaxy morphological properties do not evolve significantly for those of similar intrinsic brightness and physical size. While some previous iterations of Galaxy Zoo have involved sets of objects sharing similar, and often low, redshifts, for which this assumption approximately holds (Walmsley et al., 2022a), the Cosmic Dawn iteration’s subject set covers a much wider redshift range (primarily $0<z<1.0$, as well as many objects at higher redshift). Additionally, photometric redshift estimation from the H20 pipeline failed for approximately one fifth of the sources in the subject set. For these reasons, no redshift-based debiasing was applied to the results, with the final vote fractions remaining ‘raw’, i.e. entirely based on the visual morphology present in the images shown to volunteers. As such, classifications of objects at higher redshifts will be more adversely affected due to lower resolution, with them being more likely to be classified as non-featured (for more information on the effect of this biasing, see Walmsley et al., 2023b).
@@ 3.5.2 Galaxy Zoo Mobile gravitational lenses
Throughout the project, in addition to the ‘rare features’ question, volunteers were asked to tag any gravitational lenses they found on the Talk boards: given the rarity of such objects, we wished to explore all possible methods for compiling as complete a sample as possible, especially given that the ‘rare-features’ question is shown last to volunteers. As the project neared completion, another method was explored: in addition to the desktop version of Zooniverse, there is also a Zooniverse app for mobile devices, intended for simple and often binary ‘Yes/No’ workflows that are compatible with the mobile app’s ‘swipe left/right’ functionality. Hence, as a follow-up to the project, a workflow for finding gravitational lenses through the mobile app was launched alongside other workflows as part of a reboot of Galaxy Zoo Mobile. It consisted of 22 587 subjects, which were simply the set of galaxies that had not been fully classified in the desktop application, and removing those predicted by Zoobot to be stars, artifacts or images with ‘Bad Image Zoom’ with a 90 per cent lower bound greater than 0.5. The workflow asked volunteers if each image contained a gravitational lens or not, and used a retirement threshold of ten volunteer classifications given the far simpler, binary ‘Yes/No’ workflow. To improve speed and volunteer experience, an additional retirement rule was introduced such that a subject would be retired if it had six or more votes and every one of those votes was for the ‘no’ response.
@@ 4.1.1 Volunteer & classification number counts
GZCD ran for over six months from 21 October 2022 to 03 May 2023 and consisted of 47 347 subjects. 3 990 911 classifications were made by tens of thousands of volunteers: 74 per cent by the 10 274 logged-in volunteers along with 26 per cent from an unknown number of volunteers who made classifications without logging in. The weekly number of total classifications made by volunteers are provided in Figure 5. The peak at the end of November corresponds to an update put out on the Galaxy Zoo Talk boards, in particular highlighting the release of the H20 Image Viewer (see Section 2.4) and, correspondingly, the inclusion of image coordinates in the Galaxy Zoo subject metadata to view them beyond the Zooniverse platform. Unfortunately, the press releases (see Section 3.3) at the end of 2022 did little to increase volunteer engagement, perhaps due to a limited ability to reach would-be volunteers compared to, say, the post on the Galaxy Zoo Blog at project launch. The rise in classifications towards the end of the project was likely due to volunteers becoming more engaged as the ‘percentage complete’ bar nears completion in addition to seeing more interesting images as a result of Zoobot active learning.
@@
The first set of 16 671 subjects was fully classified by at least 40 volunteers per subject. As discussed in Section 3.4, with the introduction of Zoobot active learning to the second set, 528 of the subjects were immediately and confidently retired by Zoobot without any volunteer classifications; the remaining 30 148 each received at least one volunteer classification, of which 19 686 (65 per cent) had fewer than ten volunteer responses before being retired by Zoobot. Figure 6 shows the number of votes that each subject received from volunteers. For those with less than 40 votes, there is a skew towards low counts, due to Zoobot retiring many subjects quickly and more over time as it trains through active learning. The threshold for retirement is made obvious by the peak at 40 votes, which, in addition to subjects in the first set each receiving at least 40 votes, is amplified by 22.6 per cent of subjects in the second set also receiving at least 40 votes: for these cases, if Zoobot could not confidently classify a subject after being seen by a small number of volunteers, it was unlikely to ever retire it early, leading to those subjects hitting (or exceeding) the retirement threshold. There are many subjects that received slightly more than 40 votes, as those with just under 40 votes could receive classifications from multiple volunteers almost simultaneously before the automated retirement could be enacted.
Figure 6 also shows another peak in the distribution at higher classification counts, with 189 subjects (0.4 per cent of data set) receiving 66-109 votes each, with a median of 87 votes. Of these, 100 have randomly distributed H20 IDs: when a subject set is completed (with all subjects having reached one or more retirement criteria), Zooniverse continues to show a sample of 100 subjects, which are still classifiable but marked by a ‘Finished!’ tag, which aids in maintaining volunteer engagement before a new set becomes available. The other 89 subjects correspond to the final IDs in the first set, which all received a significant fraction of their votes ($\sim$48-93 per cent) towards the end of that set’s classification period. Why these subjects received so many classifications is uncertain, but likely due to an unknown bug with the Zooniverse platform during the course of the project runtime: they do not correlate to any particular object and are not duplicate uploads, nor are they erroneously duplicated entries in the classification results, and the vast majority of the votes come from logged-in volunteers so cannot be due to bot accounts repeatedly classifying the same subjects.